OpenStudy (anonymous):

-2m^2 + 8 / 7m^3 11m^2 - 4m + 5

OpenStudy (anonymous):

can you rewrite this please?

OpenStudy (anonymous):

use parenthesis

OpenStudy (anonymous):

or draw it out

OpenStudy (anonymous):

ok so this is the problem (-2m^2)+(8) / (7m^3)(11m62)-(4m)+5

OpenStudy (anonymous):

ups sorry its (-2m^2)+8 / (7m^3)(11m^2)-(4m)+5

OpenStudy (anonymous):

@timo86m i rewrote it

OpenStudy (anonymous):

\[-2\,{m}^{2}+{\frac {8}{7}}\, \left( m \left( 11\,{m}^{2} \right) \right) ^{-3}-4\,m+5 \] ok that last one this is what you mean?

OpenStudy (anonymous):

no its \[-2m^2+8 / 7m^3-11m^2-4m+5 = 0\]

OpenStudy (anonymous):

\[\frac{-2m^2+8} {(7m^3)(11m^2)-4m+5}\] this?

OpenStudy (anonymous):

yes exactly!

OpenStudy (anonymous):

sooo what exactly is the question?

OpenStudy (anonymous):

how to solve it

OpenStudy (anonymous):

lol sorry cherry but i have a feeling me and that other guy knew that is what you meant all along :)

OpenStudy (anonymous):

simplify it i guess

OpenStudy (anonymous):

okay so the only part i need help with is the bottom \[7m^3-11m^2-4m+5\]

OpenStudy (anonymous):

\[\frac{-2m^2 + 8 }{ 7m^3+ 11m^2 - 4m + 5}\] \[-2(x^2-4)=-2(m+2)(m-2)\] \[7m^3-11m^2-4m+5\]doesn't factor so there is nothing you can do with it

OpenStudy (anonymous):

oooh equal to zero? then solve \[(m+2)(m-2)=0\] because a fraction is only zero if the numerator is

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so what about the bottom?